Choose two words or phrases from each module, write the definitions of the two words and give examples.

((((Module 1))))
1. The Symmetric Property of Equality
2. Vertical Line Rule
3. The Commutative Property of Addition
4. The Reflexive Property of Equality
5. Slope-Intercept form of a Linear Equation
6. The Commutative Property of Multiplication
7. Order of Operations
8. Point-Slope form of a Linear Equation
9. The Associative Property of Addition
10. <=, >=, <, >
11. y-coordinate
12. The Associative Property of Multiplication
12. Combined Inequality
14. x-coordinate
15. The Identity Property of Addition
16. Relation
17. x-intercept
18. The Identity Property of Multiplication
19. Function
20. y-intercept
21. The Inverse Property of Addition (Additive Inverse)
22. Direct variation
23. Parallel Slope
24. The Inverse Property of Multiplication (Multiplicative Inverse)
25. Inverse Variation
26. Perpendicular Slope
27. The Transitive Property of Equality
28. Joint Variation
29. Cartesian Coordinate System
(((((Module 2))))
1. Factoring
2. Term
3. Greatest Common Factor
4. Like Terms
5. Difference of Squares
6. Constant
7. Sum of Squares
8. Degree of a Polynomial
9. Perfect Square Trinomial
10. Polynomial
11. Difference of Cubes
12. Monomial
13. Sum of Cubes
14. Binomial
15. Factor by Grouping
16. Trinomial
17. Quadratic Equation
18. Synthetic
19. Division Pure Quadratic Equation
Module 3
Axis of Symmetry
Perfect Square Root Discriminate
Perfect Cube Root Complete the Square
Prime Number Rational Numbers
Simplify a Radical Irrational Numbers
Rationalize a Denominator Complex Numbers
Quadratic Formula Imaginary Number
Conjugate Value of i2
Parabola
Vertex of a Parabola
Module 4 Matrix
Systems of Equations Element of a Matrix
Solving Systems Graphically Dimensions of a Matrix
Solving Systems Algebraically Column Matrix
Solving Systems by Substitution Row Matrix
Solving Systems by Addition Square Matrix
Least Common Multiple Scalar Multiplication
Systems of Inequalities Identity Matrix
Slope of a Line (m) Inverse Matrix
y-intercept of a line (b) Solving Systems by the Inverse Matrix Method
Module 5 Minor Axis of an Ellipse
Conic Section Foci of an Ellipse
Circle Equation of a Parabola (General Form)
Equation of a circle (General Form) Focus of a Parabola
Radius of a Circle Directrix of a Parabola
Center of a Circle Hyperbola
SOHCAHTOA Transverse Axis of a Hyperbola
Pythagorean Formula Conjugate Axis of a Hyperbola
Ellipse Foci of a Hyperbola
Equation of an Ellipse (General Form) Aphelion
Major Axis of an Ellipse Perihelion
Module 6 Extraneous Root
Rational Expression Remainder
Restrictions on the Domain of a Rational Expression Intermediate Value Theorem
Multiplying Rational Expressions Fundamental Theorem of Algebra
Dividing Rational Expressions Descartes' Rule of Signs
Reciprocal of a Rational Expression Even Function
Complex Fraction Odd Function
Least Common Denominator Increasing Interval
Adding Rational Expressions Decreasing Interval
Vertical Asymptote Domain
Hole Range
Horizontal Asymptote Upper & Lower Bounds
Module 7 Antilog
Radical Form of an Expression Antiln (ex)
Logarithmic Form of an Expression Product Property of Logarithms
Exponential Notation Quotient Property of Logarithms
Natural Log (ln) Power Property of Logarithms
Base of a Logarithm
Module 8 Permutations
Median

Sequence
Factorial
Mode

Arithmetic Sequence & Formula
Combinations & Formula
Quartile

Series
Probability
Deviation

Arithmetic Series & Formula
Experimental Probability
Standard Deviation

Summation & Notation
Theoretical Probability
Variance

Geometric Sequence & Formula
Random Sample
Measures of Central Tendency

Common Ratio
Stratified Sample
Box & Whisker Plot

Geometric Series & Formula
Statistics
Histogram

Fundamental Counting Principle
Mean
Stem & Leaf Plot












Answer by nabla(475)

I will do the first few modules for you here. Try doing some on your own via these examples and post the remaining modules if you still require help.
Module 1:
10. <=, >=, <, >
This is commonly used notation to make statements about the relative size or order of two quantities or objects.
"a<=b" means that a is less than or equal to b.
"a>=b" means that a is greater than or equal to b.
"a _<_ b" means that a is strictly less than b. (Take out the _--- issue with this site)
"a>b" means that a is strictly greater than b.
Further example:
(2n)(n-1)! for n>=1 can be proven easily by mathematical induction.

19. function
"The mathematical concept of a function expresses dependence between two quantities, one of which is given (the independent variable, argument of the function, or its "input") and the other produced (the dependent variable, value of the function, or "output"). A function associates a single output to each input element drawn from a fixed set, such as the real numbers." (Wikipedia)
Example:
f(x)=e^(x)sin(x)
Module 2.
5. difference of squares
These terms define themselves: a difference of squares is a difference of squared terms. Consider:
a^2-b^2
This is factored easily as a^2-b^2=(a-b)(a+b) because the non-square term is nonexistent (it is canceled).
15. factor by grouping
This phrase refers to a process of grouping terms in order to factor generally higher degreed polynomials than quadratics. Simply, determine if two pairs of terms have something the same in common. If they do, they can be factored out.
Example:
x^3-3x^2+x-3=x^3+x-3x^2-3=x(x^2+1)-3(x^2+1)=(x-3)(x^2+1)